# Networked Information Aggregation via Machine Learning

### v1

@@ C.3 Matrix Concentration for Sample Covariance
$C\cdot R_{X}^{2}\sqrt{\frac{\log d+\log(1/\delta)}{m}}<\frac{\lambda_{\min}(\Sigma)}{2}\implies m>\frac{4C^{2}R_{X}^{4}(\log d+\log(1/\delta))}{\lambda_{\min}(\Sigma)^{2}}.\qed$
@@ C.5 Bounding the Norm of Empirically Learned Predictors
$\|\beta_{k}^{emp}\|_{1}\leq\sqrt{d}\cdot\frac{\sqrt{2}Y_{\max}}{\sqrt{\lambda_{\min}(\Sigma)}}=\frac{\sqrt{2d}\cdot Y_{\max}}{\sqrt{\lambda_{\min}(\Sigma)}}.\qed$
@@ D.2 Generalization Bound for Function Compositions
Step 1: Symmetrization and High-Probability Bound. Let =$\mathcal{L}=\{(x,y)\mapsto(g(x)-y)^{2}\mid g\in\mathcal{G}_{\Lambda_{1}}(\mathcal{H})\}$ be the loss class. For any $g\in\mathcal{G}_{\Lambda_{1}}(\mathcal{H})$, the function’s output is bounded by $|g(x)|=|\sum_{h}\beta_{h}h(x)|\leq\sum_{h}|\beta_{h}||h(x)|\leq B_{h}\|\beta\|_{1}\leq\Lambda_{1}B_{h}$. The loss is therefore bounded by $(|g(x)|+|y|)^{2}\leq(\Lambda_{1}B_{h}+Y_{\max})^{2}$. Let $M_{L}=\Lambda_{1}B_{h}+Y_{\max}$.
@@
Step 2: Contraction Principle. As before, we view the loss as a composition with $\phi(z)=z^{2}$, which is $2M_{L}$-Lipschitz on $[-M_{L},M_{L}]$. Talagrand’s contraction principle gives:
@@ D.3 Bounding the Norm of Empirically Learned Predictors
$||\beta_{k}||_{1}\leq\frac{T_{k}\cdot B_{\text{basis}}\cdot Y_{\max}}{\lambda_{\min,H}}$
@@
$||\beta_{k}||_{2}\leq\left\|\left(\frac{1}{m}\mathbf{H}_{k}^{T}\mathbf{H}_{k}\right)^{-1}\right\|_{\text{op}}\left\|\frac{1}{m}\mathbf{H}_{k}^{T}y\right\|_{2}.$
@@
$||\beta_{k}||_{2}\leq\frac{\sqrt{T_{k}}B_{\text{basis}}Y_{\max}}{\lambda_{\min,H}}.$
@@
$||\beta_{k}||_{1}\leq\sqrt{T_{k}}||\beta_{k}||_{2}\leq\frac{T_{k}B_{\text{basis}}Y_{\max}}{\lambda_{\min,H}}.$

### v2

@@ C.3 Matrix Concentration for Sample Covariance
$C\cdot R_{X}^{2}\sqrt{\frac{\log d+\log(1/\delta)}{m}}<\frac{\lambda_{\min}(\Sigma)}{2}\implies m>\frac{4C^{2}R_{X}^{4}(\log d+\log(1/\delta))}{\lambda_{\min}(\Sigma)^{2}}.$
@@ C.5 Bounding the Norm of Empirically Learned Predictors
$\|\beta_{k}^{emp}\|_{1}\leq\sqrt{d}\cdot\frac{\sqrt{2}Y_{\max}}{\sqrt{\lambda_{\min}(\Sigma)}}=\frac{\sqrt{2d}\cdot Y_{\max}}{\sqrt{\lambda_{\min}(\Sigma)}}.$
@@ D.2 Generalization Bound for Function Compositions
Step 1: Symmetrization and High-Probability Bound. Let $\mathcal{L}=\{(x,y)\mapsto(g(x)-y)^{2}\mid g\in\mathcal{G}_{\Lambda_{1}}(\mathcal{H})\}$ be the loss class. For any $g\in\mathcal{G}_{\Lambda_{1}}(\mathcal{H})$, the function’s output is bounded by $|g(x)|=|\sum_{h}\beta_{h}h(x)|\leq\sum_{h}|\beta_{h}||h(x)|\leq B_{h}\|\beta\|_{1}\leq\Lambda_{1}B_{h}$. The loss is therefore bounded by $(|g(x)|+|y|)^{2}\leq(\Lambda_{1}B_{h}+Y_{\max})^{2}$. Let $M_{L}=\Lambda_{1}B_{h}+Y_{\max}$.
@@
Step 2: Contraction Principle. As before, we view the loss as a composition with $\phi(z)=z^{2}$, which is $2M_{L}$-Lipschitz on $M_{L},M_{L}]$. Talagrand’s contraction principle gives:
@@ D.3 Bounding the Norm of Empirically Learned Predictors
$\beta_{k}||_{1}\leq\frac{T_{k}\cdot B_{\text{basis}}\cdot Y_{\max}}{\lambda_{\min,H}}$
@@
$\beta_{k}||_{2}\leq\left\|\left(\frac{1}{m}\mathbf{H}_{k}^{T}\mathbf{H}_{k}\right)^{-1}\right\|_{\text{op}}\left\|\frac{1}{m}\mathbf{H}_{k}^{T}y\right\|_{2}.$
@@
$\beta_{k}||_{2}\leq\frac{\sqrt{T_{k}}B_{\text{basis}}Y_{\max}}{\lambda_{\min,H}}.$
@@
[…]